OpenStudy (afifarahman):

What's the period of Cos[(2[Pi])/3 t]+Sin[(2[Pi])/4 t]?

OpenStudy (afifarahman):

@zepdrix

OpenStudy (afifarahman):

@TheSmartOne

OpenStudy (holsteremission):

Did you mean to have \(\dfrac{2\pi}{4}\) in the sine term's argument? In any case, the period of \(\cos(ax)+\sin(bx)=\mathrm{lcm}(a,b)\).